Calculate the mass of KI in grams required to prepare 5.00 × 102 mL of a 2.3 M solution.

?g
Answer #1

From the given information first we will calculate number of moles of KI (solute) required to prepare this solution then using mole concept we will calculate mass of KI required

Given: Molarity of KI solution = 2.3 M

Volume of KI solution = 5 × 102 mL = 0.5 L

Molecular mass of KI = 166 g/mol (known)

We know

Molarity= Number of moles of solute/ volume of solution in litre

2.3 = Number of moles of KI / 0.5

Number of moles of KI= 1.15 mol

Mass of 1.15 moles of KI = number of moles × molecular mass of KI

Mass of KI = 1.15 × 166 = 190.9 g

Thus 190.9 grams of KI required to prepare this solution.